# How does the entanglement entropy of a many-body quantum system change after a single measurement?

### v3

@@ A.1 Quantum state diffusion (QSD)
$d\left|\psi(t)\right\rangle/|\psi(t)\rangle=-iHdt+\sum_{i=1}^{L}\left(\sqrt{\gamma}\left[\hat{n}_{i}-\left\langle\hat{n}_{i}\right\rangle_{t}\right]dW_{t}^{i}-\frac{\gamma}{2}\left[\hat{n}_{i}-\left\langle\hat{n}_{i}\right\rangle_{t}\right]^{2}dt\right)$ (14)
@@
where $H$ is the Hamiltonian defined in Eq. 1, $\langle\hat{n}_{i}\rangle_{t}\equiv\langle\psi(t)|\hat{n}_{i}|\psi(t)\rangle$ denotes the expectation value of the particle density at site $i$. $W_{i}^{t}$ is the Wiener process associated with Gaussian stochastic noise satisfying $dW_{t}^{i}dW_{t}^{j}=\delta_{ij}dt$. The second term on the right describes measurement effects where the monitoring strength is $\gamma$. It can be derived from the unraveling of the corresponding Lindbladian [14, 15] master equation [9]. We stress that in the QSD protocol measurements induce independent Gaussian noise at all sites and continuously in time. The time discretization $\delta t$ value is an important parameter in our numerical analysis. After numerous tests, we have found that $\delta t=0.05$ is sufficient given that the state evolution is governed by a first order equation.
@@
The state evolution equation also includes the stochastic term $dW_{t}^{i}$ which requires Ito calculus. To leading order $O(\gamma)$, we can represent the left side of Eq. 14 as the derivative of $\ln(\psi(t))$,
$d({\mathrm{l}n}(|\psi(t)\rangle))=\frac{d|\psi(t)\rangle}{|\psi(t)\rangle}-\frac{\left(d|\psi(t)\rangle\right)^{2}}{2|\psi(t)\rangle^{2}}=\frac{d|\psi(t)\rangle}{|\psi(t)\rangle}-\sum_{i=1}^{L}\left(\frac{\gamma}{2}\left[\hat{n}_{i}-\left\langle\hat{n}_{i}\right\rangle_{t}\right]^{2}\mathrm{\penalty 10000\ d}t\right)$ (15)
where, on the right, since the noise strength is proportional to $\sqrt{\gamma}$, we should consider the $(dW_{i}^{t})^{2}$ term which is of order $O(\gamma)$. As a result,
$|\psi(t+dt)\rangle=\exp\left(-iHdt+\sum_{i=1}^{L}\left[\left(\hat{n}_{i}-\langle\hat{n}_{i}\rangle_{t}\right)\delta W_{t}^{i}-\gamma\left(\hat{n}_{i}-\langle\hat{n}_{i}\rangle_{t}\right)^{2}dt\right]\right)|\psi(t)\rangle$ (16)
@@
where $\delta W_{i}^{t}$ is the Gaussian noise of zero average and standard deviation $\sqrt{\gamma\delta t}$. The state $|\psi(t)\rangle$ can also be expressed in terms of the coefficient matrix $U(t)$, $|\psi(t)\rangle=\prod_{k=1}^{N}\left[\sum_{j=1}^{L}U_{jk}(t)c_{j}^{\dagger}\right]|\text{vac}\rangle$, Eq. 2 in the main text. For any quadratic operator in the second quantization form $A=\sum_{ij}\tilde{A}_{ij}c_{i}^{\dagger}c_{j}$, it is possible to show by using the standard anti-commutation relations that $A|\psi\rangle=\prod_{k=1}^{N}\left[\sum_{j=1}^{L}\tilde{A}_{jl}U_{lk}(t)c_{j}^{\dagger}\right]|\text{vac}\rangle$, which is effectively the multiplication of the coefficients matrices $\tilde{A}U$. As a result, we can reduce the state evolution Eq. 16 to the updating of the coefficient matrix $U(t)$ to $U(t+\delta t)$ at each time step [19, 21, 56],
@@
$U(t+dt)\propto\exp\left(-i\tilde{H}dt+\delta W^{t}+(2\langle\hat{n}\rangle_{t}-\mathbf{1})\gamma dt\right)U(t),$ (17)
@@
where $\tilde{H}_{ij}=J\delta_{i\pm 1,j}$ is the $L\times L$ coefficient matrix of the Hamiltonian $H$, namely, $H=\sum_{ij}\tilde{H}_{ij}c_{i}^{\dagger}c_{j}$. $\delta W^{t}$ and $\langle\hat{n}\rangle_{t}$ are the $L\times L$ diagonal matrix whose $i$-th diagonal element equals to $\delta W^{t}_{i}$ and $\langle\hat{n}_{i}\rangle_{t}$ respectively and $\mathbf{1}$ is the $L\times L$ identity matrix.
@@
It’s worthwhile to note that, we can interpret the stochastic state evolution governed by the Ito process Eq. 14 as the one resulting from the state evolution with the following non-Hermitian Hamiltonian $H_{d}$,
$H_{d}=H+i\sum_{i=1}^{L}\hat{n}_{i}\left(\delta W_{t}^{i}+(2\langle\hat{n}_{i}\rangle-1)\gamma dt\right)$ (18)
@@
where we interpret $\delta W_{t}^{i}$ as Gaussian random numbers with no dependence on the time step $\delta t$. The low-energy dynamics of such quadratic disordered Hamilton Eq. 18 has been addressed by using the replica trick and the path-integral formalism [37, 40, 45] resulting in a Non-Linear Sigma model with symmetries that depend on both the system and the measurement protocol.
@@
1. Initialization: We initialize $U$ in the half-filling Néel state following Eq. 3, and choose as time discretization step $\delta t=0.05$.
2. State Evolution: We compute the evolution of $U(t)$ following Eq. 17, resulting in the unnormalized coefficient matrix $\tilde{U}(t+\delta t)$. Before each time step, we compute $\langle\hat{n}_{i}\rangle_{t}$, which equals to the $i$-th diagonal elements of the correlation matrix $D(t)=U(t)U(t)^{\dagger}$ defined in Eq. 6. We compute the matrix multiplication employing the fourth-order Runge-Kutta method instead of exact diagonalization because numerically the latter is much more expensive. We have checked that for $\delta t\leq 0.1$, and for all monitoring rates $\gamma$ we consider, the fourth-order Runge-Kutta method is always a good approximation.
3. Normalization: We use the QR decomposition $\tilde{U}(t+\delta t)=U(t+\delta t)R$, where $R$ is an upper-triangular matrix, to preserve the normalization condition $U^{\dagger}U=\mathbf{1}_{N}$. The procedure ensures that the new state is correctly normalized $\langle\psi(t+\delta t)|\psi(t+\delta t)\rangle=1$. This QR decomposition takes most of the computational time.
4. Compute EE: We compute the entanglement entropy and mutual information from the correlation matrix $D(t+\delta t)=U(t+\delta t)U(t+\delta t)^{\dagger}$. Note the state evolution only involves the updating of $U(t)$, so that this step does not affect the evolution, and we only choose some of time slices to compute the events we need.
@@
The QSD measurement protocol combines both measurements and Hamiltonian dynamics at each time step of the dynamics. The main advantage of this method is that the purity of the Gaussian state is conserved by construction [19]. Therefore, the time step $\delta t$ does not need to be very small to avoid non-physical errors. The procedure for updating the state matrix $U$ under the QSD protocol is given by Algorithm 2.
@@
5:  Generate Gaussian random variable $\delta W_{i}^{t}\sim N(0,\sqrt{\gamma dt})$.
@@ A.2 Quantum jump (QJ)
In the QJ protocol, the state evolves continuously by a non-Hermitian Hamiltonian until but this continuous evolution is interrupted at certain times by sudden quantum jumps physically related to a measurement, which in our case is the observation of a particle $n_{i}=1$ at site $i$. The so-called jump operator that describes this effect is defined as $L_{j}=\sqrt{\gamma}\hat{n}_{j}=\sqrt{\gamma}c_{j}^{\dagger}c_{j}$. The full dynamics including the quantum jumps can be modeled by the following equation [47, 49]:
$d|\psi(t)\rangle=-iH|\psi(t)\rangle dt+\sum_{j}\left[\frac{L_{j}|\psi(t)\rangle}{\sqrt{\langle\psi(t)|L_{j}L_{j}^{\dagger}|\psi(t)\rangle}}-|\psi(t)\rangle\right]dw_{j}(t),$ (19)
@@
3. Non-Hermitian time evolution: Evolve the state according to the effective Hamiltonian $H_{eff}=H-i\gamma/2\sum_{j=1}^{L}c_{j}^{\dagger}c_{j}$:
$\frac{d}{dt}|\psi(t)\rangle=-iH_{eff}|\psi(t)\rangle.$ (20)
4. Quantum jump: At time $t+\tau$, a quantum jump, $L_{j}=\sqrt{\gamma}\hat{n}_{j}$ representing the detection of a particle ($n_{j}=1$) at site $j$, occurs with probability $p_{j}$ given by:
$p_{j}=\frac{\gamma_{j}\langle\psi(t+\tau^{-})|L_{j}|\psi(t+\tau^{-})\rangle}{\sum_{k}\gamma_{k}\langle\psi(t+\tau^{-})|L_{k}|\psi(t+\tau^{-})\rangle}.$ (21)
5. State update: Update the state by applying the Lindblad operator $L_{j}$:
$|\psi(t+\tau^{+})\rangle\rightarrow\frac{L_{j}|\psi(t+\tau^{-})\rangle}{\sqrt{\langle\psi(t+\tau^{-})|L_{j}^{\dagger}L_{j}|\psi(t+\tau^{-})\rangle}}.$ (22)
@@ 6. Orthonormalization: Obtain the correlation matrix according to
$D_{i,i^{\prime}}=\langle\psi^{\prime}|c_{i}^{\dagger}c_{i^{\prime}}|\psi^{\prime}\rangle=\begin{cases}1,&i=i^{\prime}=j\\ 0,&(i=j,i\neq i^{\prime})\text{or}(i^{\prime}=j,i\neq i^{\prime})\\ \langle c_{i}^{\dagger}c_{i^{\prime}}\rangle-\frac{\langle c_{j}^{\dagger}c_{i^{\prime}}\rangle\langle c_{i}^{\dagger}c_{j}\rangle}{\langle c_{j}^{\dagger}c_{j}\rangle},&\text{otherwise}\end{cases}$ (23)
@@
A detailed numerical algorithm for updating the state matrix under the QJ protocol is given in Algorithm 2.
@@ A.3 Projective Measurements (PM)
This appendix introduces the PM protocol [37] which is closely related to the QJ protocol. The two main differences is that the state evolution between two measurements is governed by the Hermitian Hamiltonian and that the time between measurements $\tau$ is here a free parameter which we have decided to mimic that of the QJ protocol $\tau={\rm ln}(\eta)/(\gamma N)$ , where $\eta$ is a random value uniformly distributed in $(0,1]$, $N=L/2$ is the particle number and $\gamma$ is the measuring rate per site. At each measurement time $t+\tau$, a site $j$ is chosen randomly, and a projective measurement of the site occupation number $\hat{n}_{j}=c_{j}^{\dagger}c_{j}$ is conducted. Upon performing a measurement at time $t+\tau$ with outcome $n_{j}$, the density matrix undergoes a discontinuous change:
@@
The possible outcomes of this measurement are either $1$ or 0, reflecting the fermionic nature of the particles. Then, we can get the density-density correlation matrix $D_{ij}={\rm Tr}(\hat{\rho}\hat{n}_{i}\hat{n}_{j})-{\rm Tr}(\hat{\rho}\hat{n}_{i}){\rm Tr}(\hat{\rho}\hat{n}_{j})$ with $\hat{\rho}$ the density matrix.
Therefore, for the numerical simulation of the PM in a free fermionic system, we focus on the evolution of the correlation matrix $D(t)$. The simulation is also initialized in the Néel state $|\psi(t=0)\rangle=|10101010\ldots\rangle$, and the corresponding element of the correlation matrix can be written as $D_{ij}(t=0)=\langle\psi(t=0)|c_{i}^{\dagger}c_{j}|\psi(t=0)\rangle$. For each time interval between two PM, apply the unitary evolution $D(t+\tau)=e^{-iH\tau}D(t)e^{iH\tau}$. At time $t+\tau$, we randomly select a site $j$, and compute $p_{j}=D_{j,j}$. We then generate a random value $p_{c}$. If $p_{j}<p_{c}$, we perform the projection operation $\hat{P}_{0}(j)=1-\hat{n}_{j}$. The corresponding updated correlation matrix elements are given by [68]:
$D_{i,i^{\prime}}=\langle\psi^{\prime}|c_{i}^{\dagger}c_{i^{\prime}}|\psi^{\prime}\rangle=\begin{cases}0,&(i=j)\text{or}(i^{\prime}=j)\\ \langle c_{i}^{\dagger}c_{i^{\prime}}\rangle-\frac{\langle c_{j}^{\dagger}c_{i^{\prime}}\rangle\langle c_{i}^{\dagger}c_{j}\rangle}{1-\langle c_{j}^{\dagger}c_{j}\rangle},&\text{otherwise}\end{cases}$ (26)
@@
Following the similar calculation introduced in Appendix A.2, we apply the SVD decomposition $D=USU^{\dagger}$ to restore the Gaussian state representation where the matrix $U$ characterizes the state after the PM measurement $|\psi(t+\tau)\rangle$. We repeat the calculation until the steady state at which the EE reaches the saturation value. These steps are summarized by following the Algorithm 3.

### v4

@@ A.1 Quantum state diffusion (QSD)
$d\left|\psi(t)\right\rangle=-i\hat{H}dt|\psi(t)\rangle+\sum_{i=1}^{L}\left(\sqrt{\gamma}\left[\hat{n}_{i}-\left\langle\hat{n}_{i}\right\rangle_{t}\right]d\tilde{W}_{t}^{i}-\frac{\gamma}{2}\left[\hat{n}_{i}-\left\langle\hat{n}_{i}\right\rangle_{t}\right]^{2}dt\right)|\psi(t)\rangle\equiv\hat{K}_{t}\left|\psi(t)\right\rangle$ (15)
@@
where $\hat{H}$ is the Hamiltonian defined in Eq. (1), $\langle\hat{n}_{i}\rangle_{t}\equiv\langle\psi(t)|\hat{n}_{i}|\psi(t)\rangle$ denotes the expectation value of the particle density at site $i$. $\tilde{W}_{t}^{i}$ is the Wiener process associated with Gaussian stochastic noise satisfying $d\tilde{W}_{t}^{i}d\tilde{W}_{t}^{j}=\delta_{ij}dt$. $\hat{K}_{t}$ denotes the infinitesimal update operator collecting all $O(dt)$ and $O(\sqrt{dt})$ contributions in Eq. 15. The second term on the right describes measurement effects where the monitoring strength is $\gamma$. It can be derived from the unraveling of the corresponding Lindbladian [14, 15] master equation [9]. We stress that in the QSD protocol measurements induce independent Gaussian noise at all sites and continuously in time. The time discretization $dt$ value is an important parameter in our numerical analysis. After numerous tests, we have found that $dt=0.05$ is sufficient given that the state evolution is governed by a first order equation.
@@
The state evolution equation also includes the stochastic term $d\tilde{W}_{t}^{i}$ which requires Ito calculus. To leading order $O(\gamma dt)$, we must keep $\hat{K}_{t}^{2}$ contribution, since $(\sqrt{\gamma}d\tilde{W}_{t}^{i})^{2}=\gamma dt$,
$\left|\psi(t+dt)\right\rangle=\exp(d\hat{X}_{t})\left|\psi(t)\right\rangle\qquad d\hat{X}_{t}=\log(\mathbb{I}+\hat{K}_{t})=\hat{K}_{t}-\frac{1}{2}\hat{K}_{t}^{2}+O(dt^{\frac{3}{2}})$ (16)
where $\mathbb{I}$ is the identity operator. As a result,
$|\psi(t+dt)\rangle=\exp\left(-i\hat{H}dt+\sum_{i=1}^{L}\left[\left(\hat{n}_{i}-\langle\hat{n}_{i}\rangle_{t}\right)dW_{t}^{i}-\gamma\left(\hat{n}_{i}-\langle\hat{n}_{i}\rangle_{t}\right)^{2}dt\right]\right)|\psi(t)\rangle$ (17)
@@
where we redefine $dW_{t}^{i}=\sqrt{\gamma}d\tilde{W}_{t}^{i}$ as the Gaussian noise of zero average and standard deviation $\sqrt{\gamma dt}$. The state $|\psi(t)\rangle$ can also be expressed in terms of the coefficient matrix $U(t)$, $|\psi(t)\rangle=\prod_{k=1}^{N}\left[\sum_{j=1}^{L}U_{jk}(t)\hat{c}_{j}^{\dagger}\right]|\text{vac}\rangle$, Eq. (2) in the main text. For any quadratic operator in the second quantization form $\hat{A}=\sum_{ij}\tilde{A}_{ij}\hat{c}_{i}^{\dagger}\hat{c}_{j}$, it is possible to show by using the standard anti-commutation relations that $A|\psi\rangle=\prod_{k=1}^{N}\left[\sum_{j=1}^{L}\tilde{A}_{jl}U_{lk}(t)\hat{c}_{j}^{\dagger}\right]|\text{vac}\rangle$, which is effectively the multiplication of the coefficients matrices $\tilde{A}U$. As a result, we can reduce the state evolution Eq. (17) to the updating of the coefficient matrix $U(t)$ to $U(t+dt)$ at each time step [19, 21, 81],
@@
$U(t+dt)\propto\exp\left(-i\tilde{H}dt+dW_{t}+(2\langle\hat{n}\rangle_{t}-\mathbf{1})\gamma dt\right)U(t),$ (18)
@@
where $\tilde{H}_{ij}=J\delta_{i\pm 1,j}$ is the $L\times L$ coefficient matrix of the Hamiltonian $\hat{H}$, namely, $\hat{H}=\sum_{ij}\tilde{H}_{ij}\hat{c}_{i}^{\dagger}\hat{c}_{j}$. $dW_{t}$ and $\langle\hat{n}\rangle_{t}$ are the $L\times L$ diagonal matrix whose $i$-th diagonal element equals to $dW^{i}_{t}$ and $\langle\hat{n}_{i}\rangle_{t}$ respectively and $\mathbf{1}$ is the $L\times L$ identity matrix.
@@
It’s worthwhile to note that, we can interpret the stochastic state evolution governed by the Ito process Eq. (15) as the one resulting from the state evolution with the following non-Hermitian Hamiltonian $H_{d}$,
$\hat{H}_{d}=\hat{H}+i\sum_{i=1}^{L}\hat{n}_{i}\left(\frac{dW_{t}^{i}}{dt}+(2\langle\hat{n}_{i}\rangle-1)\gamma\right)$ (19)
@@
The low-energy dynamics of such quadratic disordered Hamilton Eq. (19) has been addressed by using the replica trick and the path-integral formalism [40, 43, 48] resulting in a Non-Linear Sigma model with symmetries that depend on both the system and the measurement protocol.
@@
1. Initialization: We initialize $U$ in the half-filling Néel state following Eq. (3), and choose as time discretization step $dt=0.05$.
2. State Evolution: We compute the evolution of $U(t)$ following Eq. (18), resulting in the unnormalized coefficient matrix $\tilde{U}(t+dt)$. Before each time step, we compute $\langle\hat{n}_{i}\rangle_{t}$, which equals to the $i$-th diagonal elements of the correlation matrix $D(t)=U(t)U^{\dagger}(t)$ defined in Eq. (6). We compute the matrix multiplication employing the fourth-order Runge-Kutta method instead of exact diagonalization because numerically the latter is much more expensive. We have checked that for $dt\leq 0.1$, and for all monitoring rates $\gamma$ we consider, the fourth-order Runge-Kutta method is always a good approximation.
3. Normalization: We use the QR decomposition $\tilde{U}(t+dt)=U(t+dt)R$, where $R$ is an upper-triangular matrix, to preserve the normalization condition $U^{\dagger}U=\mathbf{1}_{N}$. The procedure ensures that the new state is correctly normalized $\langle\psi(t+dt)|\psi(t+dt)\rangle=1$. This QR decomposition takes most of the computational time.
4. Compute EE: We compute the entanglement entropy and mutual information from the correlation matrix $D(t+dt)=U(t+dt)U^{\dagger}(t+dt)$. Note the state evolution only involves the updating of $U(t)$, so that this step does not affect the evolution, and we only choose some of time slices to compute the events we need.
@@
The QSD measurement protocol combines both measurements and Hamiltonian dynamics at each time step of the dynamics. The main advantage of this method is that the purity of the Gaussian state is conserved by construction [19]. Therefore, the time step $dt$ does not need to be very small to avoid non-physical errors. The procedure for updating the state matrix $U$ under the QSD protocol is given by Algorithm 1. The probability distribution is computed from the data contained in $10^{4}\sim 10^{5}$ quantum trajectories. The exact number of trajectories depends on the system size.
@@
5:  Generate Gaussian random variable $\delta W_{t}^{i}\sim N(0,\sqrt{\gamma dt})$.
@@ A.2 Quantum jump (QJ)
In the QJ protocol, the state evolves continuously by a non-Hermitian Hamiltonian but this continuous evolution is interrupted at certain times by sudden quantum jumps physically related to a measurement, which in our case is the observation of a particle $n_{i}=1$ at site $i$. The so-called jump operator that describes this effect is defined as $\hat{L}_{j}=\sqrt{\gamma}\hat{n}_{j}=\sqrt{\gamma}\hat{c}_{j}^{\dagger}\hat{c}_{j}$. The full dynamics including the quantum jumps can be modeled by the following equation [49, 51]:
$d|\psi(t)\rangle=-i\hat{H}|\psi(t)\rangle dt+\sum_{j}\left[\frac{\hat{L}_{j}|\psi(t)\rangle}{\sqrt{\langle\psi(t)|\hat{L}_{j}\hat{L}_{j}^{\dagger}|\psi(t)\rangle}}-|\psi(t)\rangle\right]dw_{j}(t),$ (20)
@@
3. Non-Hermitian time evolution: Evolve the state according to the effective Hamiltonian $\hat{H}_{eff}=\hat{H}-i\gamma/2\sum_{j=1}^{L}\hat{c}_{j}^{\dagger}\hat{c}_{j}$:
$\frac{d}{dt}|\psi(t)\rangle=-i\hat{H}_{eff}|\psi(t)\rangle.$ (21)
4. Quantum jump: At time $t+\tau$, a quantum jump, $\hat{L}_{j}=\sqrt{\gamma}\hat{n}_{j}$ representing the detection of a particle ($n_{j}=1$) at site $j$, occurs with probability $p_{j}$ given by:
$p_{j}=\frac{\gamma_{j}\langle\psi(t+\tau^{-})|\hat{L}_{j}|\psi(t+\tau^{-})\rangle}{\sum_{k}\gamma_{k}\langle\psi(t+\tau^{-})|\hat{L}_{k}|\psi(t+\tau^{-})\rangle}.$ (22)
5. State update: Update the state by applying the Lindblad operator $\hat{L}_{j}$:
$|\psi(t+\tau^{+})\rangle\rightarrow\frac{\hat{L}_{j}|\psi(t+\tau^{-})\rangle}{\sqrt{\langle\psi(t+\tau^{-})|\hat{L}_{j}^{\dagger}\hat{L}_{j}|\psi(t+\tau^{-})\rangle}}.$ (23)
@@ 6. Orthonormalization: Obtain the correlation matrix according to
$D_{i,i^{\prime}}=\langle\psi^{\prime}|\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}|\psi^{\prime}\rangle=\begin{cases}1,&i=i^{\prime}=j\\ 0,&(i=j,i\neq i^{\prime})\text{or}(i^{\prime}=j,i\neq i^{\prime})\\ \langle\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}\rangle-\frac{\langle\hat{c}_{j}^{\dagger}\hat{c}_{i^{\prime}}\rangle\langle\hat{c}_{i}^{\dagger}\hat{c}_{j}\rangle}{\langle\hat{c}_{j}^{\dagger}\hat{c}_{j}\rangle},&\text{otherwise}\end{cases}$ (24)
@@
A detailed numerical algorithm for updating the state matrix under the QJ protocol is given in Algorithm 2. For the quantum jump protocol, depending on the system size and the measurement strength, we obtain between $10^{3}$ and $5\times 10^{4}$ quantum trajectories . Moreover, we collect the measurements in the time interval $[0.6L,0.7L]$, when the average entanglement entropy has already reached it saturation value. As the figures show, this amount data is enough to obtain smooth distributions.
@@
5:  Generate a random value uniformly distributed $\eta\in(0,1]$ and get $\tau$ through $exp(-i\hat{H}_{eff}\tau)U(t)||=\eta$.
[…]
@@ A.3 Projective Measurements (PM)
This appendix introduces the PM protocol [40] which is closely related to the QJ protocol. The two main differences is that the state evolution between two measurements is governed by the Hermitian Hamiltonian and that the time between measurements $\tau$ is here a free parameter which we have decided to mimic that of the QJ protocol $\tau={\rm ln}(\eta)/(\gamma N)$ , where $\eta$ is a random value uniformly distributed in $(0,1]$, $N=L/2$ is the particle number and $\gamma$ is the measuring rate per site. At each measurement time $t+\tau$, a site $j$ is chosen randomly, and a projective measurement of the site occupation number $\hat{n}_{j}=\hat{c}_{j}^{\dagger}\hat{c}_{j}$ is conducted. Upon performing a measurement at time $t+\tau$ with outcome $n_{j}$, the density matrix undergoes a discontinuous change:
@@
[…]
$D_{i,i^{\prime}}=\langle\psi^{\prime}|\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}|\psi^{\prime}\rangle=\begin{cases}0,&(i=j)\text{or}(i^{\prime}=j)\\ \langle\hat{c}_{i}^{\dagger}\hat{c}_{i^{\prime}}\rangle-\frac{\langle\hat{c}_{j}^{\dagger}\hat{c}_{i^{\prime}}\rangle\langle\hat{c}_{i}^{\dagger}\hat{c}_{j}\rangle}{1-\langle\hat{c}_{j}^{\dagger}\hat{c}_{j}\rangle},&\text{otherwise}\end{cases}$ (27)
@@
Following the similar calculation introduced in Appendix A.2, we apply the SVD decomposition $D=USU^{\dagger}$ to restore the Gaussian state representation where the matrix $U$ characterizes the state after the PM measurement $|\psi(t+\tau)\rangle$. We repeat the calculation until the steady state at which the EE reaches the saturation value. These steps are summarized in Algorithm 3. We calculate from $10^{3}$ to $5\times 10^{4}$ trajectories depending on the system size and the measurement strength. The measurements are collected in the time interval $[0.6L,0.7L]$ so that the entanglement entropy has reached its saturation value. With this information, we compute the distribution of the change of entanglement entropy after each measurement.